You have $400,000 saved for retirement. Your account earns 5% interest. How much will you be able to pull out each month, if you want to be able to take withdrawals for 25 years?  
Answer #1

Ans $ 2338.36

Annuity PV Factor = P [ 1 - ( 1 + r )-n ]
        r
400000 = P* ( 1 - ((1 / (1 + 5%/12)^300)))
                             5%/12
400000 = 0.712750196003776 * P
0.00416666666666667
P = 400000 * 0.00416666666666667 / 0.712750196003776
P = 2338.36